# Spin structures and phase diagrams of the $\text{spin-}\frac{5}{2}$ triangular-lattice antiferromagnet Na2BaMn(PO4)2 under magnetic field

### v1

@@ Abstract
We combine single-crystal neutron diffraction studies and Monte Carlo simulations to determine the spin structures and finite-temperature phase diagram of the spin-5/2 triangular-lattice antiferromagnet Na2BaMn(PO4)2 in magnetic field. With the application of a magnetic field in two different directions, namely along the $c$-axis and in the $ab$-plane of the trigonal symmetry, we track the evolution of the spin structure through changes of the magnetic propagation vector. We account for these results with a minimal Heisenberg Hamiltonian that includes easy-axis anisotropy and weak, frustrated interlayer couplings in addition to intralayer exchange. Guided by representation analysis, we refine symmetry-allowed modes to the measured intensities and obtain the spin structures for all field-induced phases, which we compare quantitatively with simulated configurations. Taken together, our measurements and simulations show that frustrated interlayer exchange – rather than purely 2D physics – organizes the unexpectedly rich field-induced phases of Na2BaMn(PO4)2.
@@ I Introduction
Frustrated triangular-lattice antiferromagnets (TLAs) are considered as model systems in which exotic states can emerge. For example, the suppression of exchange interactions due to geometric frustration and a low quantum spin number ($S=1/2$) can result in strong quantum spin fluctuations at low temperatures that stabilize a quantum spin liquid state [1]. Such states exhibit several appealing properties in condensed matter physics like quasiparticle fractionalization [2], long-range entanglement [3], and topological order [4]. Novel phases of matter and interesting physical properties are not restricted only to frustrated magnet systems in the quantum limit (low spin), but can also emerge in the classical limit ($S\geq 1$[5, 6, 7] and by applying an external magnetic field [8, 9, 10, 11].
In this context, the equilateral TLAs Na2BaM(PO4)2 (with $\text{M}=\text{Co, Ni, Mn}$) are good candidate materials to explore phenomena related to quantum magnetism, since the substitution of the transition metal ion leads to drastic changes in the quantum spin number [12, 13, 14]. The ground state of the phosphate with the lowest spin ($S=1/2$), Na2BaCo(PO4)2, is proposed to form a Y-like spin supersolid phase in zero magnetic field [15, 16], and is considered an ideal material where a spin-1/2 easy-axis XXZ model can be applied [17]. In addition, the spin supersolidity is connected to a giant magnetocaloric effect which is observed during the demagnetization cooling process [15]. On the contrary, Na2BaNi(PO4)2 where $S=1$, is assumed to exhibit a spin nematic supersolid ground state that can be modeled by a spin Hamiltonian including a single-ion anisotropic term and nearest-neighbor XXZ-type exchange interactions [18]. Although this scenario is primarily supported by field-induced behavior and magnon-pair condensation at low temperatures [18], the precise nature of the zero-field ground state remains debated [19]. Na2BaMn(PO4)2 is expected to be the classical counterpart of the TLA series due to the high quantum spin number ($S=5/2$) of the Mn+2 ions [14]. Also, it is worth mentioning that all compounds within the series manifest multiple field-induced phase transitions [20, 13, 14]. Above a critical field applied along the $c$-axis, a magnetization plateau appears in a finite range of the external magnetic field where the magnetization reaches 1/3 of the saturation magnetization $M_{\text{sat}}$. Such 1/3–magnetization plateaus were first observed in GdPd2Al3 [21] and later in other TLAs such as RbFe(MoO4)2 [22] and Ba3MnNb2O9 [9]. While for Na2BaMn(PO4)2 the magnetic structure at zero field was recently investigated with neutron powder diffraction measurements [23], the spin configurations of the field-induced phases and a minimal model Hamiltonian that describes the magnetic ground state remain unknown.
@@ II.1 Experimental details
Na2BaMn(PO4)2 single crystals were grown by the high temperature flux method [12, 14] and were characterized with single crystal X-ray diffraction and specific heat measurements. Single crystal X-ray diffraction measurements were performed on a Rigaku XtaLAB Synergy-S diffractometer, using a Mo X-ray source providing a monochromatic beam with a wavelength of 0.71 Å and a two-dimensional HyPix-Arc 150 detector. The observed reflections were indexed and integrated using the data reduction program CrysalisPro [24]. The structure solution and refinement was performed with SHELX [25]. Heat capacity was measured on a Quantum Design Physical Property Measurement System (PPMS) equipped with the Heat Capacity option. Data were collected on a single crystal of Na2BaMn(PO4)2 with a mass of about 0.6 mg, employing both the semi-adiabatic time-relaxation technique and the long-pulse method described in Ref. 26.
A parallelepiped Na2BaMn(PO4)2 single crystal with mass 37.7 mg was mounted on a copper sample holder and was oriented in the (hh0)(hh0)/$(00l)$ scattering plane of the trigonal lattice. In this article, we use the hexagonal coordinate system and the momentum transfer $\bm{Q}$ is expressed in $\bm{Q}=(Q_{h},Q_{k},Q_{l})$ given in reciprocal lattice units (r.l.u.). The relation between $\bm{Q}$ and the propagation vector $\bm{k}$ is given by $\bm{Q}=\bm{G}+\bm{k}$, where $\bm{G}$ is at the $\Gamma$-point of the reciprocal lattice.
Single-crystal neutron diffraction measurements were carried out at the Institut Laue-Langevin (ILL), in Grenoble, France. Data were obtained on the two-axis thermal neutron diffractometer D23 [27]. A copper monochromator provided an unpolarized neutron beam with a wavelength of 1.282 Å. In order to reach sub-Kelvin temperatures, the sample holder was attached to the mixing chamber of a dilution stick. To investigate the evolution of the spin structure under magnetic field, measurements were performed in two geometries: with a 15 T vertical field magnet, providing magnetic field within the $ab$-plane, and with a 3.8 T horizontal field magnet, providing field along the $c$-axis. In both geometries, data at zero field were collected and combined to refine the zero-field magnetic structures at 600 and 1200 mK. The various spin structures were refined using Mag2POL software [28].
@@ II.2 Theoretical framework
For this model, the ground state spin configuration was determined through spin dynamics by solving the Landau–Lifshitz–Gilbert equation using the Spirit package [29]. The spin-wave spectrum was determined using linear spin-wave theory for noncollinear magnets [30]. Finally, Monte Carlo simulations were employed to determine the mechanical statistical properties of the system using the Vampire package [31, 32]. We used a $3\times 3\times 5$ macrocell with periodic boundaries and an equilibration time of 5000 Monte-Carlo steps, and statistical averaging also over 5000 steps.
@@ III.1 Crystal structure
Figure 1: Crystal structure of Na2BaMn(PO4)2 as determined by single crystal X-ray diffraction. View along the (a) in-plane direction, highlighting the layered structure. (b) View along the $c$-axis. (c,d) General view highlighting the connectivity of Mn ions, where (d) shows the Mn–Mn bonds used in theoretical modeling of the spin interactions, in-plane J1 (bond length is 5.375.37\,Å), out-of-plane J2 (7.0947.094\,Å), and diagonal out-of-plane Ja and Jb (both 8.9Å). Even though the bond length for Ja and Jb is the same, the bonds are not equivalent in the $P\bar{3}$ space group.
@@
All measured crystals exhibit merohedral twinning with apparent $P\bar{3}m1$ symmetry, where the mirror plane $m_{100}$ defines the twinning law. Regarding the atomic positions, the difference in assigning either space group concerns only the oxygen at the $6i$ site, where constraints of the $P\bar{3}m1$ space group give $y(\mathrm{O_{6i}})=\frac{1}{2}x(\mathrm{O_{6i}})$, while the $P\bar{3}$ space group sets the $y$ coordinate free. The refinement with the $P\bar{3}m1$ space group results in five times larger displacement parameters of $\mathrm{O_{6i}}$ and three times higher R factors than the refinement with the $P\bar{3}$ structure. Releasing the constraints and refining the structure with $P\bar{3}$ space group yields excellent quality factors and gives the twinning ratio of 70:30. The $y$ coordinate of position $\mathrm{O_{6i}}$ differs significantly from the $y=\frac{1}{2}x$ constraint ($y\approx 0.24$ vs. constrained $y\approx 0.18$) and corresponds to the rotation of the MnO6 octahedra by an angle of 9.9 from their positions in the $P\bar{3}m1$ space group. The twinning is consistent with the recently reported high temperature transition in Na2BaMn(PO4)2 [33] and explains the inconsistencies between the reports on the $P\bar{3}$ [34, 23, 35, 33] and $P\bar{3}m1$ [36, 14] space groups. It is worth mentioning that a similar twinning scheme was recently considered for Na2BaCo(PO4)2 [37].
@@ III.2 Field-induced phase transitions
Zero magnetic field heat capacity measurements on single crystals [14] and neutron powder diffraction experiments [23] revealed that long-ranged antiferromagnetic (AFM) order appears in Na2BaMn(PO4)2 at low temperatures. The system at zero field undergoes two successive magnetic transitions at $T_{N_{2}}\approx 1.28$ K (AFM2) and $T_{N_{1}}\approx 1.13$ K (AFM1) into antiferromagnetic phases that are indexed with the modulation vector $\bm{k}_{\mathrm{AFM2}}=(1/3,1/3,0.139)$ and $\bm{k}_{\mathrm{AFM1}}=(1/3,1/3,0.187)$, respectively [23]. Consistently our neutron data in Fig. 2 indicate the occurrence of spin ordering at zero field for $T<T_{N_{1}}$ with $\bm{k}_{\mathrm{AFM1}}=(1/3,1/3,0.187)$ and for $T_{N_{1}}\leq T\leq T_{N_{2}}$ with $\bm{k}_{\mathrm{AFM2}}=(1/3,1/3,0.147)$. The significant difference in $\bm{k}_{\mathrm{AFM2}}$ between our results and Ref. [23] is ascribed to the temperature dependence of the propagation vector within this phase. We also note that due to the blockage of the bulky horizontal magnet several magnetic reflections constrained in the $(hhl)$ plane are not accessible. Therefore, data could not be obtained for $0.5<Q_{l}<0.5$ r.l.u. for $\bm{H}\parallel\bm{\hat{c}}$ (see Figs 2(a),(c)).
@@
Figure 2: Color-coded intensity plot of single crystal diffraction data of Na2BaMn(PO4)2 collected at 300 and 1200 mK as a function of $\bm{Q}=(1/3,1/3,Q_{l})$ and applied magnetic field along (a), (c) the $c$-axis and (b), (d) in the $ab$-plane. White horizontal lines represent the identified phase boundaries.
@@
Based on $\chi_{\text{ac}}$ magnetic susceptibility measurements on single-crystal samples [14], when a magnetic field is applied parallel or perpendicular to the $c$-axis of the trigonal symmetry, multiple field-induced phase transitions occur, whose spin configurations are still unexplored to our knowledge. Therefore, a first essential step for determining the spin arrangement of the various spin structures is to index their magnetic propagation vector. To this aim, we performed reciprocal space scans for magnetic fields $\bm{H}\parallel\bm{\hat{c}}$ and $\bm{H}\perp\bm{\hat{c}}$, and the results are presented in Fig. 2 and Fig. 3. The AFM order is characterized by magnetic reflections arising as satellites to the nuclear Bragg reflections with modulation vector $(1/3,1/3,k_{z})$, where $k_{z}$ changes as a function of the applied field direction and magnitude.
@@
The evolution of the magnetic order at $T=300$ mK for $\bm{H}\parallel\bm{\hat{c}}$ is depicted in Fig. 2(a) where for critical fields $H>H_{c_{1}}^{\parallel}$ the system shows commensurate ordering. Three different phases can be distinguished up to the maximum investigated magnetic field of 3.8 T: (i) for $H_{c_{1}}^{\parallel}\leq H<H_{c_{2}}^{\parallel}$ the magnetic wave vector slightly shifts and becomes $(1/3,1/3,1/6)$ with a second less intense magnetic peak appearing at $(1/3,1/3,1/3)$, (ii) for $H_{c_{2}}^{\parallel}\leq H\leq H_{c_{3}}^{\parallel}$ the system enters into the 1/3 plateau phase [14] and this is accompanied by a first sudden change in $\bm{k}$ which is now located at $(1/3,1/3,0)$ [up-up-down (UUD) phase], and (iii) a second abrupt jump results to $\bm{k}=(1/3,1/3,1/3)$ for $H_{c_{3}}^{\parallel}<H\leq H_{\text{max}}^{\parallel}$. In contrast, at $T=1200$ mK for $H_{c_{1}}^{\parallel}\leq H<H_{c_{2}}^{\parallel}$, a continuous diminution of the magnetic peak intensity and a shift of $\bm{k}$ towards the K-point are observed (see Fig. 2(c)). When the system is tuned into the UUD phase, a strong enhancement of the peak intensity is detected before reaching the spin-polarized (SP) state.

### v2

@@ Abstract
We combine single-crystal neutron diffraction studies and Monte Carlo simulations to determine the spin structures and finite-temperature phase diagram of the spin-5/2 triangular-lattice antiferromagnet Na2BaMn(PO4)2 in magnetic field. With the application of a magnetic field in two different directions, namely along the $c$-axis and in the $ab$-plane of the trigonal symmetry, we track the evolution of the spin structure through changes of the magnetic propagation vector. We account for these results with a minimal Heisenberg Hamiltonian that includes easy-axis anisotropy and weak, frustrated interlayer couplings in addition to intralayer exchange. Guided by representation analysis, we refine symmetry-allowed modes to the measured intensities and obtain the spin structures for all field-induced phases, which we compare quantitatively with simulated configurations. Taken together, our measurements and simulations show that frustrated interlayer exchange – rather than purely two-dimensional (2D) physics – organizes the unexpectedly rich field-induced phases of Na2BaMn(PO4)2.
@@ I Introduction
Frustrated triangular-lattice antiferromagnets (TLAs) are considered as model systems in which exotic states can emerge. For example, the suppression of exchange interactions due to geometric frustration and a low quantum spin number ($S=1/2$) can result in strong quantum spin fluctuations at low temperatures that stabilize a quantum spin liquid state [Savary_2017]. Such states exhibit several appealing properties in condensed matter physics like quasiparticle fractionalization [Balents_2010], long-range entanglement [Broholm_2020], and topological order [Wen_1991]. Novel phases of matter and interesting physical properties are not restricted only to frustrated magnet systems in the quantum limit (low spin), but can also emerge in the classical limit ($S\geq 1$[Collins_1997, Tapp_2017, Lee_2020] and by applying an external magnetic field [Matsuda_2010, Lee_2014, Rosales_2013, Seabra_2016].
In this context, the equilateral TLAs Na2BaM(PO4)2 (with $\text{M}=\text{Co, Ni, Mn}$) are good candidate materials to explore phenomena related to quantum magnetism, since the substitution of the transition metal ion leads to drastic changes in the quantum spin number [Zhong_2019, Li_2019, Kim_2022]. The ground state of the phosphate with the lowest spin ($S=1/2$), Na2BaCo(PO4)2, is proposed to form a Y-like spin supersolid phase in zero magnetic field [Xiang_2024, Gao_2022], and is considered an ideal material where a spin-1/2 easy-axis XXZ model can be applied [Sheng_2022]. In addition, the spin supersolidity is connected to a giant magnetocaloric effect which is observed during the demagnetization cooling process [Xiang_2024]. On the contrary, Na2BaNi(PO4)2 where $S=1$, is assumed to exhibit a spin nematic supersolid ground state that can be modeled by a spin Hamiltonian including a single-ion anisotropic term and nearest-neighbor XXZ-type exchange interactions [Sheng_2025]. Although this scenario is primarily supported by field-induced behavior and magnon-pair condensation at low temperatures [Sheng_2025], the precise nature of the zero-field ground state remains debated [huang_2025u]. Na2BaMn(PO4)2 is expected to be the classical counterpart of the TLA series due to the high quantum spin number ($S=5/2$) of the Mn+2 ions [Kim_2022]. Also, it is worth mentioning that all compounds within the series manifest multiple field-induced phase transitions [Li_2020, Li_2019, Kim_2022]. Above a critical field applied along the $c$-axis, a magnetization plateau appears in a finite range of the external magnetic field where the magnetization reaches 1/3 of the saturation magnetization $M_{\text{sat}}$. Such 1/3–magnetization plateaus were first observed in GdPd2Al3 [Kitazawa_1999] and later in other TLAs such as RbFe(MoO4)2 [Svistov_2006] and Ba3MnNb2O9 [Lee_2014]. While for Na2BaMn(PO4)2 the magnetic structure at zero field was recently investigated with neutron powder diffraction measurements [Zhang_2024], the spin configurations of the field-induced phases and a minimal model Hamiltonian that describes the magnetic ground state remain unknown.
@@ II.1 Experimental details
Na2BaMn(PO4)2 single crystals were grown by the high temperature flux method [Zhong_2019, Kim_2022] and were characterized with single crystal X-ray diffraction and specific heat measurements. Single crystal X-ray diffraction measurements were performed on a Rigaku XtaLAB Synergy-S diffractometer, using a Mo X-ray source providing a monochromatic beam with a wavelength of 0.71 Å and a two-dimensional HyPix-Arc 150 detector. The observed reflections were indexed and integrated using the data reduction program CrysalisPro [Crysalis]. The structure solution and refinement was performed with SHELX [shelx]. Heat capacity was measured on a Quantum Design Physical Property Measurement System (PPMS) equipped with the Heat Capacity option. Data were collected on a single crystal of Na2BaMn(PO4)2 with a mass of about 0.6 mg, employing both the semi-adiabatic time-relaxation technique and the long-pulse method described in Ref. [Scheie_2018].
A parallelepiped Na2BaMn(PO4)2 single crystal with mass 37.7 mg was mounted on a copper sample holder and was oriented in the (hh0)(hh0)/$(00l)$ scattering plane of the trigonal lattice. In this article, we use the hexagonal coordinate system and the scattering vector $\bm{Q}$ is expressed in $\bm{Q}=(Q_{h},Q_{k},Q_{l})$ given in reciprocal lattice units (r.l.u.). The relation between $\bm{Q}$ and the propagation vector $\bm{k}$ is given by $\bm{Q}=\bm{G}+\bm{k}$, where $\bm{G}$ is at the $\Gamma$-point of the reciprocal lattice.
Single-crystal neutron diffraction measurements were carried out at the Institut Laue-Langevin (ILL), in Grenoble, France. Data were obtained on the two-axis thermal neutron diffractometer D23 [D23-LSF]. A copper monochromator provided an unpolarized neutron beam with a wavelength of 1.282 Å. In order to reach sub-Kelvin temperatures, the sample holder was attached to the mixing chamber of a dilution stick. To investigate the evolution of the spin structure under magnetic field, measurements were performed in two geometries: with a 15 T vertical field magnet, providing magnetic field within the $ab$-plane, and with a 3.8 T horizontal field magnet, providing field along the $c$-axis. In both geometries, data at zero field were collected and combined to refine the zero-field magnetic structures at 600 and 1200 mK. The various spin structures were refined using Mag2POL software [Qureshi2019].
@@ II.2 Theoretical framework
For this model, the ground state spin configuration was determined through spin dynamics by solving the Landau–Lifshitz–Gilbert equation using the Spirit package [muller_spirit:_2019]. The spin-wave spectrum was determined using linear spin-wave theory for noncollinear magnets [dos_santos_spin-resolved_2018]. Finally, Monte Carlo simulations were employed to determine the mechanical statistical properties of the system using the Vampire package [evans_atomistic_2014, vampire]. We used a $3\times 3\times 5$ macrocell with periodic boundaries and an equilibration time of 5000 Monte Carlo steps, and statistical averaging also over 5000 steps.
@@ III.1 Crystal structure
Figure 1: Crystal structure of Na2BaMn(PO4)2 as determined by single crystal X-ray diffraction. View along the (a) in-plane direction, highlighting the layered structure. (b) View along the $c$-axis. (c,d) General view highlighting the connectivity of Mn ions, where (d) shows the Mn–Mn bonds used in theoretical modeling of the spin interactions, in-plane J1 (bond length is 5.375.37\,Å), out-of-plane J2 (7.0947.094\,Å), and diagonal out-of-plane Ja and Jb (both 8.98.9\,Å). Even though the bond length for Ja and Jb is the same, the bonds are not equivalent in the $P\bar{3}$ space group.
@@
All measured crystals exhibit merohedral twinning with apparent $P\bar{3}m1$ symmetry, where the mirror plane $m_{100}$ defines the twinning law. Regarding the atomic positions, the difference in assigning either space group concerns only the oxygen at the $6i$ site, where constraints of the $P\bar{3}m1$ space group give $y(\mathrm{O_{6i}})=\frac{1}{2}x(\mathrm{O_{6i}})$, while the $P\bar{3}$ space group sets the $y$ coordinate free. The refinement with the $P\bar{3}m1$ space group results in five times larger displacement parameters of $\mathrm{O_{6i}}$ and three times higher R factors than the refinement with the $P\bar{3}$ structure. Releasing the constraints and refining the structure with $P\bar{3}$ space group yields excellent quality factors and gives the twinning ratio of 70:30. The $y$ coordinate of position $\mathrm{O_{6i}}$ differs significantly from the $y=\frac{1}{2}x$ constraint ($y\approx 0.24$ vs. constrained $y\approx 0.18$) and corresponds to the rotation of the MnO6 octahedra by an angle of 9.9 from their positions in the $P\bar{3}m1$ space group. The twinning is consistent with the recently reported high temperature transition in Na2BaMn(PO4)2 [Kajita_2024] and explains the inconsistencies between the reports on the $P\bar{3}$ [Daisuke_2014, Zhang_2024, Zhang_2024b, Kajita_2024] and $P\bar{3}m1$ [Nenert2020, Kim_2022] space groups. It is worth mentioning that a similar twinning scheme was recently considered for Na2BaCo(PO4)2 [Woodland_2025].
@@ III.2 Field-induced phase transitions
Zero magnetic field heat capacity measurements on single crystals [Kim_2022] and neutron powder diffraction experiments [Zhang_2024] revealed that long-ranged antiferromagnetic (AFM) order appears in Na2BaMn(PO4)2 at low temperatures. The system at zero field undergoes two successive magnetic transitions at $T_{N_{2}}\approx 1.28$ K (AFM2) and $T_{N_{1}}\approx 1.13$ K (AFM1) into antiferromagnetic phases that are indexed with the modulation vector $\bm{k}_{\mathrm{AFM2}}=(1/3,1/3,0.139)$ and $\bm{k}_{\mathrm{AFM1}}=(1/3,1/3,0.187)$, respectively [Zhang_2024]. Consistently our neutron data in Fig. 2 indicate the occurrence of spin ordering at zero field for $T<T_{N_{1}}$ with $\bm{k}_{\mathrm{AFM1}}=(1/3,1/3,0.187)$ and for $T_{N_{1}}\leq T\leq T_{N_{2}}$ with $\bm{k}_{\mathrm{AFM2}}=(1/3,1/3,0.147)$. The significant difference in $\bm{k}_{\mathrm{AFM2}}$ between our results and Ref. [Zhang_2024] is ascribed to the temperature dependence of the propagation vector within this phase. We also note that due to the blockage of the bulky horizontal magnet several magnetic reflections constrained in the $(hhl)$ plane are not accessible. Therefore, data could not be obtained for $0.5<Q_{l}<0.5$ r.l.u. for $\bm{H}\parallel\bm{\hat{c}}$ (see Figs 2(a),(c)).
@@
Figure 2: Color-coded intensity plots of single crystal neutron diffraction data of Na2BaMn(PO4)2 collected at 300 and 1200 mK as a function of $\bm{Q}=(1/3,1/3,Q_{l})$ and applied magnetic field along (a), (c) the $c$-axis and (b), (d) in the $ab$-plane. White horizontal lines represent the identified phase boundaries.
@@
Based on $\chi_{\text{ac}}$ magnetic susceptibility measurements on single-crystal samples [Kim_2022], when a magnetic field is applied parallel or perpendicular to the $c$-axis of the trigonal symmetry, multiple field-induced phase transitions occur, whose spin configurations are still unexplored to our knowledge. Therefore, a first essential step for determining the spin arrangement of the various spin structures is to index their magnetic propagation vector. To this aim, we performed reciprocal space scans for magnetic fields $\bm{H}\parallel\bm{\hat{c}}$ and $\bm{H}\perp\bm{\hat{c}}$, and the results are presented in Fig. 2 and Fig. 3. The AFM order is characterized by magnetic reflections arising as satellites to the nuclear Bragg reflections with modulation vector $(1/3,1/3,k_{z})$, where $k_{z}$ changes as a function of the applied field direction and magnitude.
@@
The evolution of the magnetic order at $T=300$ mK for $\bm{H}\parallel\bm{\hat{c}}$ is depicted in Fig. 2(a) where for critical fields $H>H_{c_{1}}^{\parallel}$ the system shows commensurate ordering. Three different phases can be distinguished up to the maximum investigated magnetic field of 3.8 T: (i) for $H_{c_{1}}^{\parallel}\leq H<H_{c_{2}}^{\parallel}$ the magnetic wave vector slightly shifts and becomes $(1/3,1/3,1/6)$ with a second less intense magnetic peak appearing at $(1/3,1/3,1/3)$, (ii) for $H_{c_{2}}^{\parallel}\leq H\leq H_{c_{3}}^{\parallel}$ the system enters into the 1/3 plateau phase [Kim_2022] and this is accompanied by a first sudden change in $\bm{k}$ which is now located at $(1/3,1/3,0)$ [up-up-down (UUD) phase], and (iii) a second abrupt jump results to $\bm{k}=(1/3,1/3,1/3)$ for $H_{c_{3}}^{\parallel}<H\leq H_{\text{max}}^{\parallel}$. In contrast, at $T=1200$ mK for $H_{c_{1}}^{\parallel}\leq H<H_{c_{2}}^{\parallel}$, a continuous diminution of the magnetic peak intensity and a shift of $\bm{k}$ towards the K-point are observed (see Fig. 2(c)). When the system is tuned into the UUD phase, a strong enhancement of the peak intensity is detected before reaching the spin-polarized (SP) state.